### v1

# Sensitivity Analysis for Climate Science with Generative Flow Models

Abstract
Sensitivity analysis is a cornerstone of climate science, essential for understanding phenomena ranging from storm intensity to long-term climate feedbacks. However, computing these sensitivities using traditional physical models is often prohibitively expensive in terms of both computation and development time. While modern AI-based generative models are orders of magnitude faster to evaluate, computing sensitivities with them remains a significant bottleneck. This work addresses this challenge by applying the adjoint state method for calculating gradients in generative flow models, with diffusion models as a special case. We apply this method to the cBottle generative model, an emulator of ERA5 data, to perform sensitivity analysis with respect to sea surface temperatures. Furthermore, we propose a novel gradient self-consistency check to quantitatively validate the computed sensitivities against the model’s own outputs. Our results provide initial evidence that this approach can produce reliable gradients, reducing the computational cost of sensitivity analysis from weeks on a supercomputer with a physical model to hours on a GPU, thereby simplifying a critical workflow in climate science.
1 Introduction
Sensitivity analysis is widely applied throughout climate science, e.g. for investigating the drivers of storms banomedina2025Xynthia , the importance of observations torn2008ensemble or understanding climate feedbacks bloch2024green ; van2025reanalysis . However, conducting sensitivity analysis with classical physical simulation models is prohibitively expensive, relying either on deriving and implementing adjoints by hand errico1997adjoint ; tremolet2007model ; doyle2014initial or running finite difference approximations bloch2024green . For example, a standard approach for estimating the response of the atmosphere to spatial changes in temperature fields relies on Green’s function method which requires executing an expensive General Circulation Model (GCM) for thousands of model years which can take weeks, even on a supercomputer branstator1985analysis ; holzer2000transit ; barsugli2002global ; zhou2017analyzing ; dong2019attributing ; bloch2024green ; van2025reanalysis .
Recently, there has been an explosion of development of large AI models for weather and climate. While expensive to train, they are significantly faster and differentiable by default, warranting an investigation into their potential to simplify the sensitivity analysis workflow. Modern AI models are developed in frameworks like PyTorch pytorch and JAX jax2018github , which are equipped with automatic differentiation (AD) engines baydin2018automatic . Consequently, obtaining the gradient of any scalar quantity derived from the model’s output with respect to any model input only requires a single function call to the AD engine, which has motivated their use for initial condition sensitivity studies banomedina2025Xynthia ; vonich2024predictability ; toride2025using .
Notably, flow and diffusion models sohl2015deep ; ho2020denoising ; song2021scorebased ; lipman2023flow have gained popularity in weather and climate due to their powerful generative properties and their ability to model uncertainty. Diffusion models have demonstrated state-of-the-art results in weather forecasting price2025probabilistic and are able to reproduce atmospheric states brenowitz2025climatebottlegenerativefoundation . The method presented in this paper provides a way to extract gradients from these types of models, reducing the cost of sensitivity analysis from days to weeks on a supercomputer to minutes or hours on a GPU. However, it is not as clear if these models can produce informative gradients even if they provide a good fit to the data, see Figure 1 for an illustration of this problem. To begin the work of evaluating their quality, this paper introduces a method of extracting gradients from flow models and checking their consistency with model predictions through finite differences.
Figure 1: Left. An illustration of how an overfit model (orange) has uninformative gradients when compared to a well-fit model (blue). Our experiments show that cBottle is more of the latter than the former, encouraging its use for SST sensitivity analysis. Right. A sketch of the adjoint sensitivity method in the EDM parametrization. Given the generated vector $X_{0}$ and a function to obtain the quantity of interest, $q(X_{0})$, we obtain the desired sensitivity $\frac{\partial q}{\partial c}$ by propagating $X_{t}$, the adjoint $a_{t}$ and the gradient $w_{t}$ through their respective ODEs.
2 Background and notation
Climate in a Bottle. The experiments in this work are done with the simplest Climate in a Bottle diffusion model brenowitz2025climatebottlegenerativefoundation (henceforth called cBottle). Given the day of year $\tau$, second of day $\zeta$, sea surface temperature forcing $c$, and a high dimensional random vector $\xi$, the cBottle model generates 45 atmospheric variables ($4$ 3D variables on $8$ pressure levels, 6 column variables and 7 surface variables) at a ground resolution of $\sim 100~\text{km}$. See App. C for details. cBottle has been trained on both ERA5 reanalysis data hersbach2020era5 and outputs from the ICON climate model hohenegger2023icon , providing a unique test-bed for gradient extraction methods. While the empirical results are computed with cBottle, the method presented here is not specific to cBottle and applies to generative flow models in general.
Flow models. Diffusion models sohl2015deep ; ho2020denoising ; song2021scorebased , and the more recent flow matching lipman2023flow ; liu2022flowstraightfastlearning ; albergo2023stochastic are all special cases of flow models, a family of generative models which learn a time-varying velocity field $u_{\theta}(X_{t},t)$ that transports an easy-to-sample probability distribution $p_{init}$ to a data probability distribution $p_{data}$. In what follows, we will use the parameterization of the Elucidated Diffusion Model (EDM) karras2022edm with deterministic sampling.
Conditioning. Flow models can be guided by input variables $c$ to sample from the conditional distribution $p_{data}\left(\cdot|c\right)$ zhang2024control_net ; denker2024deft ; pidstrigach2025conditioning . Therefore, the procedure for generation implies sampling $\xi$ from the standard Gaussian, picking a conditioning variable $c$ and solving the ODE
$\text{d}X_{t}=u_{\theta}(X_{t},t,c)\text{d}t,\text{ with }X_{T}=T\xi\text{ and }\xi\sim\mathcal{N}\left(0,\text{Id}\right)$
backwards in time, from $T$ to 0, obtaining the generated sample $X_{0}\sim p_{data}\left(\cdot|c\right)$.
3 Methods
Figure 2: A qualitative example for the gradient self-consistency check. Pick a quantity of interest $q$, in this case the average top of the atmosphere outgoing shortwave radiation ($I_{+}^{sw}$) in a patch of the Indian Ocean (green rectangle). Pick an SST change vector $\delta c$, add it to the current $c$, and calculate the new $q(c+\delta c)$ and the gradient $\frac{\partial q}{\partial c}(c+\delta c)$. Repeating this procedure shows that, in the $\delta c$ direction, $q$ is mostly a smooth function with sensible gradients.
We are interested in a scalar quantity derived from a generated sample $q(X_{0})$ and how this quantity varies with conditioning $\frac{\partial q}{\partial c}$. For example, in the context of cBottle, $q$ can be net outgoing radiation globally, precipitation on the coast of Peru or wind shear in the tropical Pacific, and $c$ is global sea surface temperatures. As flow models have very deep computational graphs due to the recurrent calls to the model, it is not feasible to simply use AD to obtain this derivative. We first write down a continuous counterpart of the chain-rule
$\frac{\partial q}{\partial c}=\frac{\partial q}{\partial X_{0}}\frac{\partial X_{0}}{\partial c}=\frac{\partial q}{\partial X_{0}}\int_{T}^{0}\frac{\partial X_{0}}{\partial X_{t}}\frac{\partial u_{\theta}}{\partial c}(X_{t},t,c)\text{d}t.$
The key step is to use the adjoint state $a_{t}=\frac{\partial q}{\partial X_{0}}\frac{\partial X_{0}}{\partial X_{t}}$, which satisfies the ODE
$\frac{\text{d}}{\text{d}t}a_{t}=-a_{t}\frac{\partial u_{\theta}}{\partial X_{t}}.$
Letting $w_{T-t}=a_{t}\frac{\partial X_{t}}{\partial c}$ we see the desired result is $\frac{\partial q}{\partial c}=a_{0}\frac{\partial X_{0}}{\partial c}=w_{T}$. Solving the system of ODEs
$\frac{\text{d}}{\text{d}t}\begin{bmatrix}X_{t}\\ a_{t}\\ w_{t}\end{bmatrix}=\begin{bmatrix}u_{\theta}\\ -a_{t}\frac{\partial u_{\theta}}{\partial X_{t}}\\ -a_{t}\frac{\partial u_{\theta}}{\partial c}\end{bmatrix},\text{ with }\begin{bmatrix}X_{0}\\ a_{0}\\ w_{0}\end{bmatrix}=\begin{bmatrix}X_{0}\\ \frac{\partial q}{\partial X_{0}}\\ 0\end{bmatrix}$, (1)
from 0 to $T$ yields $\frac{\partial q}{\partial c}=w_{T}$. The process is depicted in the right panel of Figure 1, and proofs can be found in pontryagin1962optimal ; chen2018neuralODEs ; schurov2022adjoint .
Gradient self-consistency check. The approximate nature of the flow model mapping from $c$ to $X_{0}$ means the reliability of its gradients is not immediately clear, see Figure 1. While the model’s outputs are typically validated for consistency against real-world data or physical models brenowitz2025climatebottlegenerativefoundation , the trustworthiness of the gradients is commonly not assessed. We bridge this gap by proposing a self-consistency check that validates the model’s gradients against its own outputs, thereby using the validation of the predictive samples to establish confidence in the model gradients. Let a finite difference in model input be $\delta c$. For small enough $\delta c$ and smooth $q$, the following approximation should hold:
$\frac{1}{2}\left(\frac{\partial q}{\partial c}(c)+\frac{\partial q}{\partial c}(c+\delta c)\right)\cdot\delta c\approx{q(c+\delta c)-q(c)}.$ (2)
Plugging these two expressions into a selected similarity metric, we can obtain a quantitative measure of the self-consistency of gradients.
ODE directional gradient propagation. While the above should confirm the correctness of the adjoint method, it is not feasible to quantitatively evaluate the gradient consistency for the entire model output, as we have to run the algorithm one scalar $q$ at a time. To do gradient self-consistency checks at scale, we use the following insight: finite difference calculations require an a priori direction $\delta c$ in which the gradients are evaluated. This motivates the use of a different method for quickly probing gradients, detailed in Appendix A.
4 Results with Climate in a Bottle
We conduct our experiments with cBottle, a diffusion model of the EDM type karras2022edm . It can be easily reparameterized as a flow model $\text{d}X_{t}=u_{\theta}(X_{t},t,c,\tau,\zeta)\text{d}t,\text{ with }X_{T}=T\xi\text{ and }\xi\sim\mathcal{N}\left(0,\text{Id}\right)$ via the probability flow ODE song2021scorebased . The conditioning variables are $c$ – sea surface temperature (SST) forcing, $\tau$ – day of the year and $\zeta$ – UTC second of day. For now, we let $c$ vary, but we keep $\tau$, $\zeta$ and $\xi$ constant.
Figure 3: Grad-check relative errors for all 45 variables generated by cBottle (see App. C.1). Lower is better.
Qualitative gradient checks. Figure 2 shows the gradient self-consistency check procedure for a selected quantity $q$, $\tau=270$ (September 27th), $\zeta=0$ and $\xi=0$. This suggests sensible gradients consistent with cBottle’s outputs are indeed produced with the adjoint sensitivity method. However, somewhat unusually for diffusion conditioning, $c$ is dependent on $\tau$: the SST forcing used by cBottle is interpolated using $\tau$ from monthly SST means (see Appendix C.2 for details on how this is done). The right-most panel in Figure 2 shows a discontinuity at $c(\tau)=c(270)$ because the model is given inconsistent inputs (a varying $c$ but fixed $\tau$). For sensitivity analyses of interannual phenomena and quantities, keeping $\tau$ and $\zeta$ fixed can make sense. However, for the analysis of seasonal phenomena, we need to consider varying $\tau$ and calculating its contribution to the gradients. See Appendix B for an extension to cover this case of the methods presented.
Scaling up gradient checks. Using the directional gradient propagation method (eq. 3) we can compute some statistics showing to what extent cBottle’s gradients are consistent with itself. As before, we fix $\xi=0$, $\zeta=0$ and $\tau=305$ (November 1st), and pick 32 pairs of conditioning SST variables exactly one day apart from October 1st to December 3rd 2021 (outside the training set of cBottle). Figure  3 shows the grad-check relative error metric for every variable by cBottle, aggregated over the 32 pairs. The grad-check relative error metrics and their uncertainties are described in Appendix A.2. The results of this specific experiment seem to suggest that variables higher in the atmosphere have better gradients. Further investigation reveals that their connection to SSTs is weaker compared to variables lower in the atmosphere, which most likely means they are a “smoother” function of SSTs. This leads to lower grad-check errors, as the approximation in equation 2 becomes more accurate. This is confirmed by ablation experiments where the finite difference $\delta c$ is reduced, see Appendix A.3.
5 Discussion and future work
Limitations. While the results in this paper are promising and suggest that cBottle has self-consistent gradients, large-scale grad-check experiments have not been conducted. Moreover, there is no guarantee that the sensitivities extracted from the model make physical sense, especially with out-of-distribution conditioning, like SSTs expected due to climate change (see Appendix C.2).
Future work. Studies on the physical correctness of gradients must be performed, both with the presented algorithm and its extension to seasonal sensitivity analysis (B). To get a representative value of any sensitivity and its uncertainties, we need to compute the mean and variance of $\frac{\partial q}{\partial c}$ over the random vector $\xi$. The gradients will need to be further averaged over different times of day $\zeta$ to include different weather states across the globe, and perhaps neighboring values of SSTs $c$ and day of year $\tau$, for a fuller picture of the gradients’ behavior in a certain month or season.
A promising avenue for greatly extending the applications of this method is model guidance. To condition the model to generate desired weather states from the exact posterior (for example, show a realistic hurricane in the Atlantic), we can train a separate model zhang2024control_net ; denker2024deft ; pidstrigach2025conditioning that uses a guiding variable $y=G(X_{0})$, where $G$ is an observation operator. Gradients with respect to $y$ can then be pulled through the small guiding model. This method would allow calculations of gradients of anything with respect to anything, including variables not included in the original model, for example CO2 concentrations, aerosol optical depths or ocean salinity.
Appendix A ODE directional gradient propagation
Solving the ODE system
$\frac{\text{d}}{\text{d}t}\begin{bmatrix}X_{t}\\ v_{t}\end{bmatrix}=\begin{bmatrix}u_{\theta}(X_{t},t,c)\\ \frac{\partial u_{\theta}}{\partial X_{t}}(X_{t},t,c)v_{t}+\frac{\partial u_{\theta}}{\partial c}(X_{t},t,c)\delta c\end{bmatrix},\text{ with }\begin{bmatrix}X_{T}\\ v_{T}\end{bmatrix}=\begin{bmatrix}T\xi\\ 0\end{bmatrix}$. (3)
from $T$ to 0 yields the desired directional derivative $v_{0}=\frac{\partial X_{0}}{\partial c}\delta c$ in the $\delta c$ direction. Choosing the L1–norm as a measure of model’s gradient consistency, we can write down the grad-check metric $GC$:
$GC(X_{0},c,\delta c)=\left\lVert\frac{1}{2}\left(\frac{\partial X_{0}}{\partial c}(c)\delta c+\frac{\partial X_{0}}{\partial c}(c+\delta c)\delta c\right)-\left(X_{0}(c+\delta c)-X_{0}(c)\right)\right\rVert_{1}.$ (4)
A.1 Proof
Starting from the ODE
$\frac{\text{d}}{\text{d}t}X_{t}=u_{\theta}(X_{t},t,c)$
take the partial derivative $\frac{\partial}{\partial c}$ while keeping $\theta$ and $t$ constant
$\frac{\text{d}}{\text{d}t}\left(\frac{\partial X_{t}}{\partial c}\right)_{\theta,t}=\left(\frac{\partial u_{\theta}}{\partial c}\right)_{\theta,t}=\left(\frac{\partial u_{\theta}}{\partial X_{t}}\right)_{\theta,t,c}\left(\frac{\partial X_{t}}{\partial c}\right)_{\theta,t}+\left(\frac{\partial u_{\theta}}{\partial c}\right)_{\theta,t,X_{t}},$
where we have used the fact that the derivative operators $\frac{\text{d}}{\text{d}t}$ and $(\partial_{c})_{\theta,t}$ commute. Letting $h_{t}=\left(\frac{\partial X_{t}}{\partial c}\right)_{\theta,t}$ and neglecting to explicitly show what is kept constant in situations where everything is, we get
$\frac{\text{d}}{\text{d}t}h_{t}=\frac{\partial u_{\theta}}{\partial X_{t}}h_{t}+\frac{\partial u_{\theta}}{\partial c}.$
Instead of calculating the full Jacobians $\frac{\partial u_{\theta}}{\partial X_{t}}$ and $\frac{\partial u_{\theta}}{\partial c}$, which quickly becomes impossible in high-dimensional situations like image generation, we instead multiply the above equation by $\delta c$ from the right and with another change of variables $v_{t}=h_{t}\delta c$, we obtain the desired ODE:
$\frac{\text{d}}{\text{d}t}v_{t}=\frac{\partial u_{\theta}}{\partial X_{t}}v_{t}+\frac{\partial u_{\theta}}{\partial c}\delta c.$
This method leverages the efficient Jacobian Vector Product algorithms of PyTorch or JAX to directly obtain $\frac{\partial u_{\theta}}{\partial X_{t}}v_{t}+\frac{\partial u_{\theta}}{\partial c}\delta c$ — this can be done at the additional cost of just one forward pass. Finally, it is clear that $v_{T}=\frac{\partial X_{T}}{\partial c}\delta c=0$ as at the starting point the velocity field guided by $c$ has not yet pushed $X_{T}$ at all. Putting this together with the forward pass in the same ODE system yields equation 3.
Full derivative. We can also extend this to get the full derivative (see Appendix B):
$\frac{\text{d}}{\text{d}t}\left(\frac{\partial X_{t}}{\partial c}\right)_{\theta,t}=\left(\frac{\partial u_{\theta}}{\partial X_{t}}\right)_{\theta,t,c,\tau}\left(\frac{\partial X_{t}}{\partial c}\right)_{\theta,t}+\left(\frac{\partial u_{\theta}}{\partial c}\right)_{\theta,t,X_{t},\tau}+\left(\frac{\partial u_{\theta}}{\partial\tau}\right)_{\theta,t,X_{t},c}\left(\frac{\partial\tau}{\partial c}\right)_{\theta,t},$
which, multiplied from the right by $\delta c$, neglecting to explicitly write down what is kept constant when everything is, and using the expression for $\frac{\partial\tau}{\partial c}$ from Appendix B, yields
$\frac{\text{d}}{\text{d}t}v_{t}=\frac{\partial u_{\theta}}{\partial X_{t}}v_{t}+\frac{\partial u_{\theta}}{\partial c}\delta c+\frac{\partial u_{\theta}}{\partial\tau}\frac{\tau_{i+1}-\tau_{i}}{c(\tau_{i+1})-c(\tau_{i})}\delta c.$
A.2 Grad-check relative error metric
With the directional gradient method, we can form the grad-check metric for all generated variables at the same time. While this is a good way of probing the gradients of the whole model output, it does not guarantee self-consistency for any $q$, as $q(X_{0})$ might be a highly non-smooth function of $X_{0}$.
Let $y=\Omega(X_{0})$, a variable generated by cBottle (for example, outgoing long-wave radiation globally), where $\Omega$ is a simple slicing operator. The grad-check relative error metric for $y$ is defined as
$\text{RGC}=\frac{\mathbb{E}_{\xi,c,\delta c}\left\lVert\frac{1}{2}\left(\frac{\partial y}{\partial c}(c)\delta c+\frac{\partial y}{\partial c}(c+\delta c)\delta c\right)-(y(c+\delta c)-y(c))\right\rVert_{1}}{\mathbb{E}_{\xi,c,\delta c}\left\lVert y(c+\delta c)-y(c)\right\rVert_{1}}.$
Let’s unpack this. The numerator is simply the two terms of equation 2 plugged into an L1–norm. The expectation is taken over the random vector $\xi$, to probe gradients of more and less likely weather states given the conditioning — although for a given pair of points, $\xi$ should be the same for both, otherwise we would have to use a probabilistic metric like CRPS. We also take the expectation over the pair $c$ and $c+\delta c$ to probe gradients at multiple points on the conditioning manifold and in different directions $\delta c$. In the case of cBottle and for an even more comprehensive metric, we should average over $\tau$ and $\zeta$ too.
The denominator acts as a normalization term, which allows comparison of the metric across different variables $y$. Not only does it make the metric unitless, but also counteracts the effects of some variables simply changing less than others on huge swaths of the globe — for example, in cBottle, sea ice fraction differences are always very small compared to other variables.
It is worth investigating new versions of this metric in future work, but the principles guiding the design of the current one seem sensible and should be reused.
We will now explain how the uncertainties in the RGC were calculated (the error-bars in Figure 3). As the RGC is a quotient of two random variables, we used the approximation
$\operatorname{Var}\!\left(\tfrac{X}{Y}\right)=\frac{1}{n\,\mathbb{E}[Y]^{2}}\left(\operatorname{Var}(X)+\left(\tfrac{\mathbb{E}[X]}{\mathbb{E}[Y]}\right)^{2}\operatorname{Var}(Y)-2\tfrac{\mathbb{E}[X]}{\mathbb{E}[Y]}\,\operatorname{Cov}(X,Y)\right),$
where $X$ and $Y$ are two random variables and $n$ is the number of samples used for the approximation.
For the “Overall” error, we have combined the standard deviations (square root of the above variance) in quadrature, as if the variables are independent, except for the profile variables, where high correlation was apparent — we just reduced the weight of their RGC standard deviation values in the quadrature combination.
A.3 Ablations
See Figure 4.
[…]
(c) Same as (b), but we have used the seasonal sensitivity algorithm (the full derivative), as described in Appendix B.
Figure 4: As mentioned in the main text, the trend of variables in the upper atmosphere seemingly having worse gradients disappears once we decrease the step-size $\delta c$, compare (a) to (b). This also highlights the limitations of the grad-check relative error metric (A.2), as some variables that change very little in large parts of the globe (like sea ice fraction and SST) tend to have higher relative errors due to normalization. In (c) we also show the performance of the full derivative algorithm, which is slightly worse than the control (b), most likely due to discretization errors — this requires further investigation.
Appendix B Getting full derivatives with the adjoint sensitivity method
In the case in which the inputs to the flow model are dependent on each other, as is the case with cBottle’s forcing SSTs $c$ and day of year $\tau$, we might be interested in total derivatives, in addition to the partial derivatives obtainable through the adjoint sensitivity method. To reiterate, obtaining partial derivatives (from a velocity-reparameterized EDM) with the adjoint sensitivity method implies solving the system of ODEs
$\frac{\text{d}}{\text{d}t}\begin{bmatrix}X_{t}\\ a_{t}\\ w_{t}\\ v_{t}\end{bmatrix}=\begin{bmatrix}u_{\theta}\\ -a_{t}\frac{\partial u_{\theta}}{\partial X_{t}}\\ -a_{t}\frac{\partial u_{\theta}}{\partial c}\\ -a_{t}\frac{\partial u_{\theta}}{\partial\tau}\end{bmatrix},\text{ with }\begin{bmatrix}X_{0}\\ a_{0}\\ w_{0}\\ v_{0}\end{bmatrix}=\begin{bmatrix}X_{0}\\ \frac{\partial q}{\partial X_{0}}\\ 0\\ 0\end{bmatrix}$
from $t=0$ to $t=T$, where $w_{T}=\frac{\partial q}{\partial c}$ and $v_{T}=\frac{\partial q}{\partial\tau}$. In practice, calculating the additional derivative $\frac{\partial u_{\theta}}{\partial\tau}$ adds negligible computational cost, as the backpropagation process through the computational graph happens the same way and with the same starting tangent $a_{t}$. The total derivative is then
$\frac{\text{d}q}{\text{d}c}=\frac{\partial q}{\partial c}+\frac{\partial q}{\partial\tau}\frac{\partial\tau}{\partial c}=w_{T}+v_{T}\frac{\partial\tau}{\partial c}.$
In the case of cBottle, we know that $c(\tau)$ is linearly interpolated between the nearest two mid-month monthly mean SSTs (see Appendix C.2)
$c(\tau)=c(\tau_{i})+\frac{\tau-\tau_{i}}{\tau_{i+1}-\tau_{i}}\left(c(\tau_{i+1})-c(\tau_{i})\right),$
and so
$\frac{\partial\tau}{\partial c}=\left(\frac{\partial c}{\partial\tau}\right)^{-1}=\frac{\tau_{i+1}-\tau_{i}}{c(\tau_{i+1})-c(\tau_{i})}.$
Otherwise, a more general method of approximating $\frac{\partial\tau}{\partial c}$ is to train a regressor and to get this gradient with an AD call.
Appendix C cBottle details
C.1 Generated variables
See Table 1.
Short name Description Units
Profile variables @ 1000, 850, 700, 500, 300, 200, 50, 10 $\mathrm{h}\mathrm{P}\mathrm{a}$
T air temperature (profile) $\mathrm{K}$
U zonal wind (profile) $\mathrm{m}\,\mathrm{s}^{-1}$
V meridional wind (profile) $\mathrm{m}\,\mathrm{s}^{-1}$
Z geopotential height $\mathrm{m}$
Column variables
clivi column integrated cloud ice $\mathrm{k}\mathrm{g}\,\mathrm{m}^{-2}$
cllvi column integrated cloud water $\mathrm{k}\mathrm{g}\,\mathrm{m}^{-2}$
rlut TOA outgoing longwave radiation $\mathrm{W}\,\mathrm{m}^{-2}$
rsut TOA outgoing shortwave radiation $\mathrm{W}\,\mathrm{m}^{-2}$
rsds surface downwelling shortwave radiation $\mathrm{W}\,\mathrm{m}^{-2}$
tcwv column integrated water vapour $\mathrm{k}\mathrm{g}\,\mathrm{m}^{-2}$
Surface variables
pr precipitation flux $\mathrm{k}\mathrm{g}\,\mathrm{m}^{-2}\,\mathrm{s}^{-1}$
pres_msl mean sea level pressure $\mathrm{P}\mathrm{a}$
sic sea ice concentration (fractional)
sst sea surface temperature $\mathrm{K}$
tas 2m air temperature $\mathrm{K}$
uas zonal wind at 10m $\mathrm{m}\,\mathrm{s}^{-1}$
vas meridional wind at 10m $\mathrm{m}\,\mathrm{s}^{-1}$
Table 1: Variables grouped into profile, column, and surface categories with their ICON names, descriptions, and units.
C.2 The nature of SST conditioning data
As mentioned in the results section, cBottle is using one of its conditioning variables, the day of year $\tau$, to obtain the sea surface temperature forcing $c(\tau)$. This procedure is mentioned but not detailed in their paper, and so we will describe here. They use the AMIP dataset input4MIPs.CMIP6Plus.CMIP.PCMDI.PCMDI-AMIP-1-1-9.ocean.mon.tosbcs.g which can be downloaded with a script from here. This dataset contains the estimated global mid-month, monthly mean SST from January 1870 to December 2022, all dated on the 16th of the month (except February, which has the 15th). If the model is given a day of year $\tau$, $c(\tau)$ is linearly interpolated from the nearest two data points:
$c(\tau)=c(\tau_{i})+\frac{\tau-\tau_{i}}{\tau_{i+1}-\tau_{i}}\left(c(\tau_{i+1})-c(\tau_{i})\right),$
where $\tau_{i}$ is the mid-month day-of-year of month $i$ and $\tau$ satisfies $\tau_{i}\leq\tau<\tau_{i+1}<\tau+31$.
For simplicity, we have maintained that $\tau$ is the day of the year, but it would be more precise to say it is the day of the year plus the day fraction: $\tau=1.5$ means January 1st 12:00 UTC.
Figure 5: A 3D slice of the SST manifold used to train cBottle, the years 1991 to 2021 are shown. Because of the high density of interpolated values (here denoted by thin grey lines), we expect gradients obtained in an SST change direction corresponding to seasonal SST changes to be sensible. The same cannot be said for directions pointing away from the manifold, which most of them will be.